OpenStudy (anonymous):

Find an equation of the line. Write the equation using function notation. Through (2, -20) and (-8, 30)

OpenStudy (anonymous):

\[y-y_{1}= (\frac{y _{2}-y _{1} }{ x _{2}-x _{1}})(x-x _{1})\] y+20=m(x-2) \[m=\frac{ 30+20 }{ -8-2 }=-5\] so y+20=-5x+10 5x+y+10=0

OpenStudy (anonymous):

thx!